# Towards the Worst-case Robustness of Large Language Models

### v3

@@ 4 Lower Bounding Worst-case Robustness
In this section, we aim to provide a theoretical lower bound for the worst-case robustness of randomized defenses, defined as $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. We begin by discussing the formulation of randomized smoothing-based certified robustness in Sec. 4.1. Next, in Sec. 4.2 and Sec. 4.3, we show that the certified robustness of any smoothed function $g$ can be solved using a greedy algorithm from the fractional knapsack solver when $f$ is a bounded function, and this bound can be improved using dynamic programming from the 0-1 knapsack solver when $f$ is a binary function.
@@ 4.1 Formulation of Certified Robustness for LLMs
Inspired by Cohen et al. [23], we define certified robustness using randomized smoothing as the following functional optimization problem:
@@
Given a base model $f:\mathcal{X}\to\mathbb{R}$ and a smoothing distribution $p(\bm{z}|\bm{x})$, we define the smoothed function $g:\mathcal{X}\to\mathbb{R}$ as $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. Let $g(\bm{x})=p_{A}$, assume $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$ for some distance metric $\mathcal{D}$. We define randomized smoothing as the optimization for finding minimal output $g(\bm{x}_{adv})$ over all functions (or the worst) $f$ and $\bm{x}_{adv}$:
$p_{adv}=\min_{f,\bm{x}_{adv}}g(\bm{x}_{adv})=\min_{f,\bm{x}_{adv}}\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$ (1)
[…]
@@
If $p_{adv}\geq\tau$ for a given threshold $\tau$, we say the function $g$ is considered certifiably robust for input $\bm{x}$ within range $d$.
We discuss how to apply this technique to common tasks like classification and detection in Sec. B.6. In the following, we show three ways to apply this technique to certify the safety of language models.
@@
Way I: Certifying the detector. Let $\mathcal{V}$ be the vocabulary, $N$ be the sequence length. The base detector $f:\mathcal{V}^{N}\to[0,1]$ outputs values close to $1$ if the input is harmful and close to 00 if it is not. The user specifies the threshold $\tau$ to adjust the conservativeness of the detector. If we can show that, for a given base detector and $g(\bm{x})=p_{A}$, $g(\bm{x}_{adv})$ remains greater than $\tau$ for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$, then the detector $g(\bm{x})$ is certifiably robust within the distance $d$.
@@
Way II: Certifying “sure”. Most current jailbreaking attacks force the model to output “sure” as the first word. If we can certify that the model does not output “sure,” we can provably defend against these attacks. Here, $f:\mathcal{V}^{N}\to[0,1]$ represents the probability that the base language model outputs “sure,” and the threshold is set as $\tau=\frac{1}{|\mathcal{V}|}$. If we can show that $g(\bm{x}_{adv})$ is smaller than $\tau$ for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$, then the detector $g(\bm{x})$ is successfully certified within $d$. However, this approach is not applicable to attacks where the attackers do not set the optimization target to “sure.”
@@
Way III: Certifying the Output of a Language Model. Given a language model $f:\mathcal{V}^{N}\to\mathcal{V}^{N}$ and a judgment oracle $\mathcal{O}:\mathcal{V}^{N}\to\{0,1\}$, we construct a smoothed function $g(\bm{x})=\mathbb{E}[\mathcal{O}(f(\bm{z}))]$ (i.e., returning 1 when the output is safe and 0 when unsafe), which represents the probability that $f(\bm{z})$ produces a benign output. If we can show that $g(\bm{x}_{adv})$ is greater than $\tau$, this demonstrates that the output of the language model is safe with at least probability $\tau$. This definition is general, as the judgment oracle can encompass other benchmarks, enabling certification of various desired properties (e.g., coding, math, CoT, safety). However, although we obtain a tight lower bound for Eq. 1, we may still be unable to derive a practical bound for this definition. This limitation may be addressed in the future by incorporating additional neural network-dependent constraints. See Sec. I.1 for details.
@@
Therefore, in the main paper, we focus exclusively on certifying a safety detector.
@@ 4.2 Certified Robustness on Bounded $f$
Previous researchers have addressed certified robustness for simple distributions, such as Gaussian distributions [23], masking distributions (with a fixed masking ratio) [29], and synonym distributions [28]. However, these methods are not applicable to a general distribution. To address this, we propose a solution for solving the constrained optimization problem in Eq. 1 for any smoothing distribution. We begin by reviewing the Fractional Knapsack Problem:
@@
(The Revised Fractional Knapsack Problem). Given a set of items, each item $\bm{z}$ has a weight $p(\bm{z}|\bm{x})$ and a value $-p(\bm{z}|\bm{x}_{adv})$. The goal is to select items (or fractions of items) such that the total weight $\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})$ must be strictly equal to the knapsack’s capacity $p_{A}$, while maximizing the total value $-\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$, where $f(\bm{z})\in[0,1]$ denotes the fraction of each item chosen.
There are two differences between Definition 4.2 and the traditional fractional knapsack problem. First, Definition 4.2 is a minimization problem rather than a maximization problem, but this can be addressed by defining the value of each item as $-p(\bm{z}|\bm{x}_{adv})$ instead of $p(\bm{z}|\bm{x}_{adv})$. Second, Definition 4.2 requires that the total weight of items must be strictly equal to the knapsack’s capacity $p_{A}$, rather than less than or equal to it. Since the greedy algorithm of fractional knapsack solvers always finds a solution that precisely fits the knapsack (as demonstrated in Algorithm 1), this constraint does not pose an issue.
The solution to the Fractional Knapsack Problem relies on a well-known greedy algorithm: we prioritize items by their value-to-weight ratio $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$, selecting items in descending order of this ratio until the capacity $p_{A}$ is reached. This approach is optimal because it maximizes the contribution of each item (or fraction of an item) per unit weight added to the knapsack [50, 51].
Therefore, to solve Definition 4.2, we simply enumerate all possible $\bm{z}$, sort them by $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ in descending order, and select items until the cumulative weight reaches $p_{A}$, as shown in Algorithm 1. Each time we select a $\bm{z}$, we consume $p(\bm{z}|\bm{x})$ from $p_{A}$, but add $p(\bm{z}|\bm{x}_{adv})$ to $p_{adv}$. Consequently, we refer to the negative value-to-weight ratio $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ as the trading rate. The larger the trading rate, the greater the increase in $p_{adv}$, the better “our trade” is.
Algorithm 1 Fractal Knapsack Solver for (1)
@@
1:  Sort $\bm{z}\in\mathcal{X}$ by $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ (descending), and initialize $W,V\leftarrow 0$.
2:  for each $\bm{z}$ in sorted order do
3:     if $W+p(\bm{z}|\bm{x})\leq p_{A}$ then
4:        $W\leftarrow W+p(\bm{z}|\bm{x})$, $V\leftarrow V+p(\bm{z}|\bm{x}_{adv})$.
5:     else
6:        Select fraction of $\bm{z}$ to fill remaining $p_{A}-W$ by $V\leftarrow V+\left(p(\bm{z}|\bm{x}_{adv})\cdot\frac{p_{A}-W}{p(\bm{z}|\bm{x})}\right)$
7:     end if
8:  end for
@@
(Proof in Sec. C.1 and [50]). Algorithm 1 exactly solves the functional minimization part in Eq. 1.
After solving the functional minimization $\min_{f}$, one can simply enumerate all $\bm{x}_{adv}$ and run the knapsack solver for each $\bm{x}_{adv}$, identifying the minimal one for solving the input minimization problem $\min_{\bm{x}_{adv}}$. However, it is usually not necessary to enumerate all $\bm{x}_{adv}$ due to an equivalence: if this $f$ performs worst on a given $\bm{x}_{adv}$, then there exists another $f$ that performs worst on a different $\bm{x}_{adv}$. In other words, the functional minimization $\min_{f}$ already implicitly solves most of the input minimization $\min_{\bm{x}_{adv}}$.
@@ 4.3 Certified Robustness on Binary $f$
Note that the tightness of Algorithm 1 relies on the assumption that the hypothesis set of $f$ includes all functions mapping $\mathcal{V}^{N}$ to $[0,1]$. If we restrict the hypothesis set to functions that map to $\{0,1\}$ (i.e., hard functions that output only 0 or 1), this reduces to a 0-1 Knapsack problem, yielding a tighter result.
@@
(The Revised 0-1 Knapsack Problem). Given a set of items, for each item $\bm{z}$, it has a weight $p(\bm{z}|\bm{x})$ and a value $-p(\bm{z}|\bm{x}_{adv})$. The goal is to select items such that the total weight $\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})$ must be strictly equal to the knapsack’s capacity $p_{A}$, while maximizing the total value $-\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$, where $f(\bm{z})\in\{0,1\}$ indicates whether each item is chosen.
There are still two differences between Definition 4.4 and the traditional 0-1 knapsack problem. First, the minimization problem can still be converted to a maximization problem by defining the value of each item as $-p(\bm{z}|\bm{x}_{adv})$ instead of $p(\bm{z}|\bm{x}_{adv})$. Second, the requirement that the total weight must be strictly equal to the knapsack’s capacity $p_{A}$, rather than less than or equal to it, introduces additional complexity. While the traditional 0-1 knapsack problem can be reduced to this problem by introducing a slack variable, this problem cannot be reduced to the traditional 0-1 knapsack problem (as it requires an additional constraint). In other words, this problem is more challenging than the traditional 0-1 knapsack problem. Fortunately, we can still devise a dynamic programming approach to solve it; see Sec. C.2 for details.
Remark 4.5.
Note that the certified robustness obtained by these two solvers is indeed tight, i.e., there exists a worst-case bounded function $f$ for the fractional knapsack solver and a worst-case binary function $f$ for the 0-1 knapsack solver that satisfy all constraints in Eq. 1, with $g(\bm{x}_{adv})$ equal to the lower bound obtained by our solvers. In other words, the bound for Eq. 1 cannot be further improved. In the future, one might modify Eq. 1 to introduce additional constraints on the base model $f$ (e.g., Lipschitz continuity [52, 53]) to obtain a tighter bound.

### v4

@@ 4 Lower Bounding Worst-case Robustness
In this section, we aim to provide a theoretical lower bound for the worst-case robustness of randomized defenses, defined as $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. We begin by discussing the formulation of randomized smoothing-based certified robustness in Sec.˜4.1. Next, in Sec.˜4.2 and Sec.˜4.3, we show that the certified robustness of any smoothed function $g$ can be solved using a greedy algorithm from the fractional knapsack solver when $f$ is a bounded function, and this bound can be improved using dynamic programming from the 0-1 knapsack solver when $f$ is a binary function.
@@ 4.1 Formulation of Certified Robustness for LLMs
Given a base model $f:\mathcal{X}\to\mathbb{R}$ and a smoothing distribution $p(\bm{z}|\bm{x})$, we define the smoothed function $g:\mathcal{X}\to\mathbb{R}$ as $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. Let $g(\bm{x})=p_{A}$ and assume $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$ for some distance metric $\mathcal{D}$. We define the certification problem as finding the minimal output of $g(\bm{x}_{adv})$ over all possible $\bm{x}_{adv}$:
$p_{adv}:=\min_{\bm{x}_{adv}}g(\bm{x}_{adv})=\min_{\bm{x}_{adv}}\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv}),\quad\text{s.t.\ \ }\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d.$ (1)
@@
If $p_{adv}\geq\tau$ for a given threshold $\tau$, we say the function $g$ is certifiably robust for input $\bm{x}$ within distance $d$.
As far as we know, this definition encompasses all application scenarios of randomized smoothing. For example, in image classification Cohen et al. (2019); Salman et al. (2019), $g$ represents the smoothed probability of the correct class, and $\tau$ is set to 0.5 (i.e., the probability of the correct class should exceed 0.5). The goal is to find a worst-case $\bm{x}_{adv}$ within $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$ that minimizes $g(\bm{x}_{adv})$. If $g(\bm{x}_{adv})$ remains greater than $\tau=0.5$, the smoothed function $g$ is considered certifiably robust within distance $d$. See Sec.˜B.6 for additional application scenarios. In the following, we discuss three ways to apply this technique to certify the safety of LLM.
@@
Way I: Certifying the detector. Let $\mathcal{V}$ be the vocabulary, $N$ be the sequence length. The base detector $f:\mathcal{V}^{N}\to[0,1]$ outputs values close to $1$ if the input is harmful and close to 0 if it is not. The user specifies the threshold $\tau$ to adjust the conservativeness of the detector. If we can show that, for a given base detector and $g(\bm{x})=p_{A}$, $g(\bm{x}_{adv})$ remains greater than $\tau$ for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$, then the detector $g(\bm{x})$ is certifiably robust within the distance $d$.
@@
Way II: Certifying “sure”. Most current jailbreaking attacks force the model to output “sure” as the first word Zou et al. (2023). If we can certify that the model does not output “sure”, we can provably defend against these attacks. Here, $f:\mathcal{V}^{N}\to[0,1]$ represents the the probability that the base language model does not outputs “sure”, and the threshold is set as $\tau=1-\frac{1}{|\mathcal{V}|}$. If we can show that $g(\bm{x}_{adv})$ is still larger than $\tau$ for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$, then the detector $g(\bm{x})$ is successfully certified within $d$. However, this approach is not applicable to attacks where the attackers do not set the optimization target to “sure”.
@@
Way III: Certifying the Output of an LLM. Given a language model $f:\mathcal{V}^{N}\to\mathcal{V}^{N}$ and a judgment oracle $\mathcal{O}:\mathcal{V}^{N}\to\{0,1\}$, we construct a smoothed function $g(\bm{x})=\mathbb{E}[\mathcal{O}(f(\bm{z}))]$ (i.e., returning 1 when the output is safe and 0 when unsafe), which represents the probability that $f(\bm{z})$ produces a benign output. If we can show that $g(\bm{x}_{adv})$ is greater than $\tau$, this demonstrates that the output of $f$ is safe with at least probability $\tau$. This definition is general, as the judgment oracle can encompass other benchmarks, enabling certification of various desired properties (e.g., coding, math, CoT, safety). However, although we obtain a tight lower bound for Eq.˜1 in Sec.˜4.2, we may still be unable to derive a practical bound for this definition. This limitation may be addressed in the future by incorporating additional neural network-dependent constraints. See Sec.˜I.1 for details.
@@
Therefore, in the main paper, we focus exclusively on certifying a safety detector (i.e., Way I).
@@ 4.2 Certified Robustness on Bounded $f$
Previous researchers have addressed certified robustness for simple distributions, such as Gaussian distributions Cohen et al. (2019), masking distributions (with a fixed masking ratio) Zeng et al. (2023), and synonym distributions Ye et al. (2020). However, these methods are not applicable to a general distribution. To address this, we propose a solution for solving the constrained optimization problem in Eq.˜1 for any smoothing distribution.
We regard randomized smoothing as a technique for obtaining a lower bound on $g(\bm{x}_{adv})$ by relaxing the problem of finding the worst-case output of a given smoothed function $f$ to any smoothed $f^{\prime}$:
$\min_{\bm{x}_{adv}}g(\bm{x}_{adv})\geq\min_{\bm{x}_{adv}}\min_{f^{\prime}\in\mathcal{F}}\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x}_{adv}),\;\text{s.t.\ }\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x})=p_{A},\ \mathcal{D}(\bm{x},\bm{x}_{adv})\leq d,$ (2)
where $\mathcal{F}=\{f^{\prime}\mid f^{\prime}:\mathcal{X}\to[0,1]\}$ when $f$ is a bounded function ($\mathcal{X}\to[0,1]$) [^2] Without loss of generality, any bounded function can be normalized into this range., and $\mathcal{F}=\{f^{\prime}\mid f^{\prime}:\mathcal{X}\to\{0,1\}\}$ when $f$ is a binary function ($\mathcal{X}\to\{0,1\}$). To obtain this lower bound, we will show that the functional optimization $\min_{f^{\prime}\in\mathcal{F}}$ is similar to a fractional knapsack problem when $f^{\prime}$ is a bounded function, and to a 0-1 knapsack problem when $f^{\prime}$ is a binary function. For the case of bounded functions, we begin by establishing the equivalence between the functional minimization and the following knapsack problem:
@@
(The Revised Fractional Knapsack Problem). Given a set of items, each item $\bm{z}$ has a weight $p(\bm{z}|\bm{x})$ and a value $p(\bm{z}|\bm{x}_{adv})$. The goal is to select fractions of items such that the total weight $\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x})$ must be strictly equal to the knapsack’s capacity $p_{A}$, while minimizing the total value $\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x}_{adv})$, where $f^{\prime}(\bm{z})\in[0,1]$ denotes the fraction of each item chosen.
There are two differences between Definition˜4.2 and the traditional fractional knapsack problem. First, Definition˜4.2 is a minimization problem rather than a maximization problem, but they are equivalent by defining the item value as $p(\bm{z}|\bm{x}_{adv})$ instead of $p(\bm{z}|\bm{x}_{adv})$. Second, Definition˜4.2 requires that the total weight of items must be strictly equal to the knapsack’s capacity $p_{A}$, rather than less than or equal to it. Since the greedy algorithm of fractional knapsack solvers always finds a solution that precisely fits the knapsack (as shown in Algorithm˜1), this constraint is not an issue.
The solution to the Fractional Knapsack Problem relies on a well-known greedy algorithm: prioritizing items by value-to-weight ratio $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$, selecting items in descending order of this ratio until the capacity $p_{A}$ is reached. This approach is optimal because it maximizes the contribution of each item per unit weight added to the knapsack Aho & Hopcroft (1974); Cormen et al. (2022).
Therefore, to solve Definition˜4.2, we can simply enumerate all possible $\bm{z}$, sort them by $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ in descending order, and select items until the cumulative weight reaches $p_{A}$, as shown in Algorithm˜1. Each time we select a $\bm{z}$, we consume $p(\bm{z}|\bm{x})$ from $p_{A}$, but add $p(\bm{z}|\bm{x}_{adv})$ to $p_{adv}$. Consequently, we refer to the negative value-to-weight ratio $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ as the trading rate. The larger the trading rate, the greater the increase in $p_{adv}$, the better “our trade” is.
Algorithm 1 Fractional Knapsack Solver for equation 1
@@
1: Sort $\bm{z}\in\mathcal{X}$ by $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ (descending), and initialize $W,V\leftarrow 0$.
2:For each $\bm{z}$ in sorted order:
[…]
4:  else:  Select fraction of $\bm{z}$ to fill remaining $p_{A}-W$ by $V\leftarrow V+\left(p(\bm{z}|\bm{x}_{adv})\cdot\frac{p_{A}-W}{p(\bm{z}|\bm{x})}\right)$
@@
(Proof in Sec.˜C.1 and Aho & Hopcroft (1974)). Algorithm˜1 exactly solves the functional minimization part in Eq.˜2.
Solving the input minimization $\min_{\bm{x}_{adv}}$. After solving the functional minimization $\min_{f}$, solving the input minimization $\min_{\bm{x}_{adv}}$ is typically much simpler. This is because the relaxation in Eq.˜2 typically introduces symmetrization with respect to $\bm{x}_{adv}$. Intuitively, for any $\bm{x}_{adv}$, the worst-case $f^{\prime}$ corresponding to this $\bm{x}_{adv}$ performs equivalently. If a given $f^{\prime}$ performs worst on a specific $\bm{x}_{adv}$, there exists another $f^{\prime\prime}$ that performs worst on a different $\bm{x}_{adv}$. For example, in $\ell_{2}$ settings for image classification, given an $\bm{x}_{adv}$ satisfying $\|\bm{x}_{adv}-\bm{x}\|_{2}=d$, the worst-case $f^{\prime}$ is a linear classifier with a decision boundary orthogonal to the line from $\bm{x}_{adv}$ to $\bm{x}$ when smoothing distribution is isotropic Gaussian distribution. Regardless of the choice of $\bm{x}_{adv}$, the worst-case $f^{\prime}$ is always such a linear classifier, resulting in the same $g(\bm{x}_{adv})$. Similarly, in our work, for any $\bm{x}_{adv}$ such that $\|\bm{x}_{adv}-\bm{x}\|_{0}=d$, these $\bm{x}_{adv}$ values consistently yield items with the same weight, value, and value-to-weight ratio, leading the knapsack program to produce identical results (See Sec.˜D.6 for the formal construction of this equivalence). In conclusion, we view randomized smoothing as relaxing the function $f$ to the hypothesis class $\mathcal{F}$, introducing symmetrization so that we only need to solve $\min_{f^{\prime}\in\mathcal{F}}$ rather than $\min_{\bm{x}_{adv}}$.
Tightness of the bound. For the case where $f:\mathcal{X}\to[0,1]$ is a bounded function, we make a tightness claim similar to Cohen et al. (2019): If $g(\bm{x})=p_{A}$ is the only known information about $f$, it is impossible to certify a higher $g(\bm{x}_{adv})$ than the output of the knapsack solver for Eq.˜2. This is because the knapsack algorithm constructs an $f^{\prime}$ such that $\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x})=p_{A}$, where $f^{\prime}$ is defined by the selection of each item as the function output. If $g(\bm{x})=p_{A}$ is the only known information about $f$, then $f$ could be $f^{\prime}$, meaning that $\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})$ cannot exceed the knapsack solver output $\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x})$. Thus, our bound is black-box tight, i.e., by only knowing one point information $g(\bm{x})=p_{A}$, there indeed exists a worst-case $f^{\prime}$ such that this bound holds.
Equivalence to previous results. Note that the result of relaxing Definition˜4.1 via Eq.˜2 and solving with fractional knapsack solvers is equivalent to prior randomized smoothing results Cohen et al. (2019); Teng et al. (2020); Ye et al. (2020). On one hand, these bounds are all black-box tight (in the sense that $g(\bm{x})=p_{A}$ is the only known information about $f$), so they must be identical. On the other hand, we provide a formal proof of this equivalence for Gaussian and laplace distributions in Sec.˜D.5. This equivalence bridges our knapsack-based approach with established randomized smoothing frameworks, reinforcing the robustness of our theoretical findings.
@@ 4.3 Certified Robustness on Binary $f$
Note that the tightness of Algorithm˜1 relies on the assumption that the hypothesis set of $f$ includes all functions mapping $\mathcal{X}$ to $[0,1]$. If we restrict the hypothesis set to functions that map to $\{0,1\}$ (i.e., hard functions that output 0 or 1), this reduces to a 0-1 Knapsack problem, yielding a tighter result.
@@
(The Revised 0-1 Knapsack Problem). Given a set of items, for each item $\bm{z}$, it has a weight $p(\bm{z}|\bm{x})$ and a value $p(\bm{z}|\bm{x}_{adv})$. The goal is to select items such that the total weight $\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x})$ must be strictly equal to the knapsack’s capacity $p_{A}$, while minimizing the total value $\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x}_{adv})$, where $f^{\prime}(\bm{z})\in\{0,1\}$ indicates whether each item is chosen.
There are still two differences between Definition˜4.4 and the traditional 0-1 knapsack problem. First, the minimization problem can still be converted to a maximization problem by defining the value of each item as $p(\bm{z}|\bm{x}_{adv})$ instead of $p(\bm{z}|\bm{x}_{adv})$. Second, the requirement that the total weight must be strictly equal to the knapsack’s capacity $p_{A}$, rather than less than or equal to it, introduces additional complexity. While the traditional 0-1 knapsack problem can be reduced to this problem by introducing a slack variable, this problem cannot be reduced to the traditional 0-1 knapsack problem (as it requires an additional constraint). In other words, this problem is more challenging than the traditional 0-1 knapsack problem. Fortunately, we can still devise a dynamic programming approach to solve it; see Sec.˜C.2 for details.
Tightness of the bound. Note that this bound is strictly better than those obtained by fractional knapsack solvers. This is because the hypothesis set of bounded functions includes binary functions, allowing the worst-case function in fractional knapsack solvers to be selected as a binary function in this section. Additionally, this bound is also black-box tight (if $g(\bm{x})=p_{A}$ and $f:\mathcal{X}\to\{0,1\}$ are the only known information about $f$). In other words, the bound for Definition˜4.1 cannot be further improved without additional information. In the future, one might modify Definition˜4.1 to introduce further constraints on the base model $f$ (e.g., Lipschitz continuity Chen et al. (2024a); Delattre et al. (2024)) to achieve a tighter bound.